i am older than my sister, 15 years ago i was 3 times my sister age, 10 years ago i was 2 times my sister age 15 years ago.
what is my age now?

Answer by stanbon(75887)

i am older than my sister, 15 years ago i was 3 times my sister age, 10 years ago i was 2 times my sister age 15 years ago.
what is my age now?
--------
Equations:
b-15 = 3(s-15)
b-10 = 2(s-10)
--------------------
Modify;:
b - 3s = -30
b - 2s = -10
------------------
Subtract ans solve for "s"::
s = 20 (sister's age now)
----
Solve for "b":
b - 3s = -30
b - 60 = -30
b = 30 (brother's age now)
---------
Cheers,
Stan H.
-----------